# Evaluating Self-Supervised Learning in Medical Imaging: A Benchmark for Robustness, Generalizability, and Multi-Domain Impact

### v1

@@ D.1.2 Linear Evaluation with Limited Labels
Figure  15 illustrates the mean test accuracy of ViT-Small across datasets under 1%, 10%, and 100% label availability, comparing models initialized with ImageNet1K weights to those with random initialization (hatched bars).
@@ D.2.1 Effect of SSL Method
Figures 16, 18, and 17 present the density distributions of AUROC scores for various SSL methods across different model backbones and initialization schemes. These visualizations provide insights into the performance of each method in OOD detection under distinct configurations. Notably, MoCo v3 consistently achieves the highest AUROC scores with ViT-Small backbones, both when randomly initialized and when initialized with ImageNet1K weights. In contrast, NNCLR demonstrates superior OOD detection performance when paired with a ResNet-50 backbone. Figure 16 suggests that for SimCLR models, ImageNet1K initialization with a ResNet-50 backbone results in a distinctively low AUROC score for OOD detection. This variability in performance highlights the impact of both the SSL method and the model architecture on OOD detection capabilities, emphasizing the importance of selecting the right combination for optimal results.
@@ D.2.2 Effect of Backbone Architecture
The choice of backbone architecture is crucial for OOD detection performance. Figure 19 demonstrates how the choice of backbone affects the distribution of OOD AUROC scores for each method, strongly suggesting that ViT-Small is the superior option among all SSL methods in general. A comparison between Figures 5 and 17 for random initialization setting reveals that ViT-Small architectures exhibit a lower variance and a higher density of AUROC scores in the upper range, indicating their effectiveness in OOD detection. Specifically, as depicted in Figure 17, ViT-Small models, particularly those employing MoCo v3 and SimCLR, show pronounced peak densities near AUROC scores of $0.96$. This is significantly higher compared to the ResNet-50 models shown in Figure 5, where the AUROC scores are more broadly distributed with a maximum around $0.93$. Furthermore, this observation holds true for the ImageNet1K initialization setting as well. Figures  16 and 18 demonstrate that the ViT-Small architecture consistently exhibits a higher density of large AUROC scores compared to the ResNet-50 architecture.
To investigate the impact of backbone architecture in more detail, we analyze the effect of backbone among different $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs in Figures 20 and 21. For random initialization, we observe an improvement in OOD scores when ViT-Small is used across many of the $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs. In contrast, Figures 22 and 23 reveal a shift in backbone preferences when models are initialized with ImageNet1K weights. Notably, SimCLR, DINO, ReSSL, and MoCo v3 tend to favor ViT-Small for the majority of $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs, whereas BYOL, NNCLR, and VICReg demonstrate a preference for ResNet-50. For Barlow Twins, there is no clear preference for either backbone.
@@ D.2.3 Effect of Initialization
For the OOD detection task, we compare the importance of ImageNet1K-supervised weight initialization with that of random initialization. Figures 24 and 25 illustrate the effect of initialization across ResNet-50 based models and $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs. It is evident that the effect of initialization depends on both the choice of SSL model and the specific $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pair. For many datasets, methods such as SimCLR, ReSSL, MoCo v3, and VICReg tend to prefer random initialization over ImageNet1K initialization, whereas BYOL and DINO favor ImageNet1K initialization. Additionally, models trained on smaller datasets like BreastMNIST and RetinaMNIST specifically benefit from ImageNet1K initialization across all SSL methods for OOD detection. This observation indicates that, for the ResNet-50 backbone, smaller datasets gain significant advantages from supervised ImageNet1K weights.
@@
Next, we investigate whether this observation also holds for the ViT-Small backbone. Figures 26 and 27 illustrate the effect of initialization for ViT-Small architectures. Compared to ResNet-50, it can be observed that the effect of initialization is diminished, as the differences between AUROC scores are closer to zero for many $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs. Despite this, several models still favor random initialization over ImageNet1K initialization for when trained with datasets such as OCTMNIST, OrganAMNIST, OrganCMNIST, and OrganSMNIST. BYOL, DINO, and ReSSL methods trained with BreastMNIST prefers ImageNet1K initialization for better OOD detection. However, unlike the ResNet-50 backbone, this preference does not extend consistently across other SSL methods. These findings suggest that while ViT-Small may be less sensitive to initialization strategies overall, certain models and datasets still benefit from appropriate weight initialization.
@@ D.3 Generalizability
In Figures  29 and  30, we provide a detailed analysis of cross-dataset transfer performance across the MedMNIST collection for various self-supervised learning methods.
@@
Comparison of ImageNet1K vs. Random Initialization for In-Domain and Cross-Dataset Performance: In Table 12, we provide the results of paired $t$-tests to analyze the effect of ImageNet1K initialization on the in-domain performance and generalizability of the learned representations. To evaluate the impact of ImageNet1K initialization compared to random initialization, we conducted paired $t$-tests for each test dataset across all cross-dataset training combinations including the in-domain setting where test and train splits come from the same dataset. The results demonstrate statistically significant improvements $(p<0.05)$ across all datasets, with varying degrees of benefit. OCTMNIST showed the largest relative improvement ($22.08\pm 31.16\%$), followed by PathMNIST ($16.95\pm 32.10\%$) and OrganSMNIST ($13.55\pm 29.62\%$). Even datasets with higher baseline accuracies benefited from ImageNet1K initialization, with BloodMNIST improving from 80.23 to 88.41 ($13.52\pm 27.58\%$ relative gain). The smallest improvements were observed in BreastMNIST ($3.47\pm 6.47\%$), DermaMNIST ($3.76\pm 4.48\%$), PneumoniaMNIST ($3.93\pm 7.28\%$) and RetinaMNIST ($4.00\pm 7.56\%$), though these gains remained statistically significant. Notably, all improvements were achieved with strong statistical significance $(p<0.05)$, providing robust evidence for the benefit of ImageNet1K initialization in self-supervised pretraining across diverse medical imaging domains.
@@
BloodMNIST $80.52$ $88.41$ $13.52\pm 27.58\%$ $0.0000$
BreastMNIST $79.68$ $82.26$ $3.47\pm 6.47\%$ $0.0000$
DermaMNIST $70.67$ $73.26$ $3.76\pm 4.48\%$ $0.0000$
OctMNIST $58.33$ $68.24$ $22.08\pm 31.16\%$ $0.0000$
OrganAMNIST $81.54$ $88.94$ $12.79\pm 26.96\%$ $0.0000$
OrganCMNIST $79.57$ $85.15$ $10.86\pm 27.45\%$ $0.0002$
OrganCMNIST $66.31$ $72.30$ $13.55\pm 29.62\%$ $0.0000$
PathMNIST $72.90$ $81.86$ $16.95\pm 32.10\%$ $0.0000$
PneumoniaMNIST $83.50$ $86.52$ $3.93\pm 7.28\%$ $0.0000$
RetinaMNIST $48.82$ $50.63$ $4.00\pm 7.56\%$ $0.0000$
TissueMNIST $52.99$ $57.46$ $9.50\pm 13.30\%$ $0.0000$
@@ E Conclusion & Future Work
Comparison of Augmentations
For consistency in our comparisons, we maintain the default augmentation strategies specified in the solo-learn library for each SSL method. Future research could explore the impact of domain-specific augmentation strategies on model performance, providing insights into the relationship between data augmentation choices and in-domain performance, OOD detection, and generalizability.
Extensive Hyperparameter Search
We conducted a grid search across learning rates and weight decay parameters during the encoder pre-training phase and then linear evaluation. Future research directions could explore a broader hyperparameter space by investigating the effects of varying batch sizes, temperature scaling factors, and momentum coefficients in the optimization process.
Additional Multi-Domain Datasets
We evaluated multi-domain learning across two distinct dataset combinations: one comprising same-modality datasets (Organ{A,C,S}) and another incorporating datasets from different imaging modalities. Future research could extend this multi-domain investigation by systematically exploring a broader spectrum of domain combinations, encompassing both intra-modality (datasets within the same imaging modality) and cross-modality scenarios. This expanded analysis would provide deeper insights into the scalability and limitations of multi-domain learning approaches when handling varying numbers of different domains and diverse modality combinations.
Architectural Variations
Our investigation focused on ResNet-50 and ViT-Small architectures as representative examples of convolutional and transformer-based approaches. Future studies could explore the impact of architecture using different variations of convolutional and transformer-based backbones. Additionally, investigating the role of architectural components like attention mechanisms, skip connections, and normalization layers specifically in the medical imaging context could reveal design principles better suited for healthcare applications.

### v2

@@ D.1.2 Linear Evaluation with Limited Labels
Figure  14 illustrates the mean test accuracy of ViT-Small across datasets under 1%, 10%, and 100% label availability, comparing models initialized with ImageNet1K weights to those with random initialization (hatched bars).
@@ D.2.1 Effect of SSL Method
Figures 15, 17, and 16 present the density distributions of AUROC scores for various SSL methods across different model backbones and initialization schemes. These visualizations provide insights into the performance of each method in OOD detection under distinct configurations. Notably, MoCo v3 consistently achieves the highest AUROC scores with ViT-Small backbones, both when randomly initialized and when initialized with ImageNet1K weights. In contrast, NNCLR demonstrates superior OOD detection performance when paired with a ResNet-50 backbone. Figure 15 suggests that for SimCLR models, ImageNet1K initialization with a ResNet-50 backbone results in a distinctively low AUROC score for OOD detection. This variability in performance highlights the impact of both the SSL method and the model architecture on OOD detection capabilities, emphasizing the importance of selecting the right combination for optimal results.
@@ D.2.2 Effect of Backbone Architecture
The choice of backbone architecture is crucial for OOD detection performance. Figure 18 demonstrates how the choice of backbone affects the distribution of OOD AUROC scores for each method, strongly suggesting that ViT-Small is the superior option among all SSL methods in general. A comparison between Figures 5 and 16 for random initialization setting reveals that ViT-Small architectures exhibit a lower variance and a higher density of AUROC scores in the upper range, indicating their effectiveness in OOD detection. Specifically, as depicted in Figure 16, ViT-Small models, particularly those employing MoCo v3 and SimCLR, show pronounced peak densities near AUROC scores of 0.960.960.96. This is significantly higher compared to the ResNet-50 models shown in Figure 5, where the AUROC scores are more broadly distributed with a maximum around 0.930.930.93. Furthermore, this observation holds true for the ImageNet1K initialization setting as well. Figures  15 and 17 demonstrate that the ViT-Small architecture consistently exhibits a higher density of large AUROC scores compared to the ResNet-50 architecture.
To investigate the impact of backbone architecture in more detail, we analyze the effect of backbone among different $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs in Figures 19 and 20. For random initialization, we observe an improvement in OOD scores when ViT-Small is used across many of the $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs. In contrast, Figures 21 and 22 reveal a shift in backbone preferences when models are initialized with ImageNet1K weights. Notably, SimCLR, DINO, ReSSL, and MoCo v3 tend to favor ViT-Small for the majority of $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs, whereas BYOL, NNCLR, and VICReg demonstrate a preference for ResNet-50. For Barlow Twins, there is no clear preference for either backbone.
@@ D.2.3 Effect of Initialization
For the OOD detection task, we compare the importance of ImageNet1K-supervised weight initialization with that of random initialization. Figures 23 and 24 illustrate the effect of initialization across ResNet-50 based models and $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs. It is evident that the effect of initialization depends on both the choice of SSL model and the specific $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pair. For many datasets, methods such as SimCLR, ReSSL, MoCo v3, and VICReg tend to prefer random initialization over ImageNet1K initialization, whereas BYOL and DINO favor ImageNet1K initialization. Additionally, models trained on smaller datasets like BreastMNIST and RetinaMNIST specifically benefit from ImageNet1K initialization across all SSL methods for OOD detection. This observation indicates that, for the ResNet-50 backbone, smaller datasets gain significant advantages from supervised ImageNet1K weights.
@@
Next, we investigate whether this observation also holds for the ViT-Small backbone. Figures 25 and 26 illustrate the effect of initialization for ViT-Small architectures. Compared to ResNet-50, it can be observed that the effect of initialization is diminished, as the differences between AUROC scores are closer to zero for many $(\mathcal{P}_{\text{ID}},\mathcal{P}_{\text{OOD}})$ pairs. Despite this, several models still favor random initialization over ImageNet1K initialization for when trained with datasets such as OCTMNIST, OrganAMNIST, OrganCMNIST, and OrganSMNIST. BYOL, DINO, and ReSSL methods trained with BreastMNIST prefers ImageNet1K initialization for better OOD detection. However, unlike the ResNet-50 backbone, this preference does not extend consistently across other SSL methods. These findings suggest that while ViT-Small may be less sensitive to initialization strategies overall, certain models and datasets still benefit from appropriate weight initialization.
Table 12: AUROC Differences (self-supervised ImageNet1K minus supervised ImageNet1K) for ResNet-50 models. Values are reported as $\Delta$AUC $\pm$ 95% CI. Significant differences ($p<0.05$) are highlighted in green if self-supervised weights performs better and red if supervised weights performs better. If the absolute difference is smaller than 0.050.050.05, they are highlighted in light green and light red instead.
Dataset MoCo v3 DINO BYOL SimCLR
BloodMNIST 0.003±0.004-0.003$\pm$ 0.004- 0.003 ± 0.004 $\pm 0.005$ 0.005 0.008±0.0160.008$\pm$ 0.0160.008 ± 0.016 0.057±0.048-0.057$\pm$ 0.048- 0.057 ± 0.048
BreastMNIST 0.033±0.043-0.033$\pm$ 0.043- 0.033 ± 0.043 0.398±0.166-0.398$\pm$ 0.166- 0.398 ± 0.166 0.001±0.0030.001$\pm$ 0.0030.001 ± 0.003 0.066±0.068-0.066$\pm$ 0.068- 0.066 ± 0.068
Dermamnist 0.095±0.066-0.095$\pm$ 0.066- 0.095 ± 0.066 0.010±0.0070.010$\pm$ 0.0070.010 ± 0.007 0.008±0.013-0.008$\pm$ 0.013- 0.008 ± 0.013 0.091±0.1130.091$\pm$ 0.1130.091 ± 0.113
OctMNIST 0.047±0.0350.047$\pm$ 0.0350.047 ± 0.035 0.048±0.059-0.048$\pm$ 0.059- 0.048 ± 0.059 0.004±0.005-0.004$\pm$ 0.005- 0.004 ± 0.005 0.099±0.1840.099$\pm$ 0.1840.099 ± 0.184
OrganAMNIST 0.047±0.0750.047$\pm$ 0.0750.047 ± 0.075 0.040±0.0910.040$\pm$ 0.0910.040 ± 0.091 0.072±0.098-0.072$\pm$ 0.098- 0.072 ± 0.098 0.073±0.1240.073$\pm$ 0.1240.073 ± 0.124
OrganCMNIST 0.026±0.096-0.026$\pm$ 0.096- 0.026 ± 0.096 0.157±0.1150.157$\pm$ 0.1150.157 ± 0.115 0.147±0.112-0.147$\pm$ 0.112- 0.147 ± 0.112 0.078±0.149-0.078$\pm$ 0.149- 0.078 ± 0.149
OrganSMNIST 0.062±0.174-0.062$\pm$ 0.174- 0.062 ± 0.174 0.131±0.1580.131$\pm$ 0.1580.131 ± 0.158 0.009±0.072-0.009$\pm$ 0.072- 0.009 ± 0.072 0.104±0.203-0.104$\pm$ 0.203- 0.104 ± 0.203
PathMNIST 0.022±0.020-0.022$\pm$ 0.020- 0.022 ± 0.020 0.270±0.080-0.270$\pm$ 0.080- 0.270 ± 0.080 0.027±0.015-0.027$\pm$ 0.015- 0.027 ± 0.015 0.032±0.141-0.032$\pm$ 0.141- 0.032 ± 0.141
PneumoniaMNIST 0.025±0.020-0.025$\pm$ 0.020- 0.025 ± 0.020 0.011±0.0050.011$\pm$ 0.0050.011 ± 0.005 0.029±0.0130.029$\pm$ 0.0130.029 ± 0.013 0.302±0.1420.302$\pm$ 0.1420.302 ± 0.142
RetinaMNIST $\pm 0.002$ 0.002 0.069±0.054-0.069$\pm$ 0.054- 0.069 ± 0.054 0.031±0.040-0.031$\pm$ 0.040- 0.031 ± 0.040 0.036±0.164-0.036$\pm$ 0.164- 0.036 ± 0.164
TissueMNIST 0.018±0.021-0.018$\pm$ 0.021- 0.018 ± 0.021 0.347±0.2410.347$\pm$ 0.2410.347 ± 0.241 0.001±0.002-0.001$\pm$ 0.002- 0.001 ± 0.002 0.007±0.014-0.007$\pm$ 0.014- 0.007 ± 0.014
Table 12 summarizes the AUROC differences—computed as the AUROC of models initialized with self-supervised ImageNet1K weights minus that of models initialized with supervised ImageNet1K weights—across various datasets for four SSL methods, using an $\alpha$ level of 0.05. As detailed in the table, the magnitude and direction of the differences vary considerably with the SSL method. For instance, DINO shows a strong preference for initializing with supervised ImageNet1K weights for BreastMNIST, while it favors self-supervised ImageNet1K initialization for TissueMNIST. Similarly, SimCLR indicates a marked preference for self-supervised ImageNet1K initialization in PneumoniaMNIST, but a preference for supervised ImageNet1K weights in BloodMNIST and BreastMNIST. In contrast, MoCo v3 and BYOL tend to yield more modest differences overall, suggesting less pronounced preferences. These results highlight that the choice of self-supervised initialization significantly influences OOD detection performance in a dataset-dependent manner.
@@ D.3 Generalizability
In Figures  28 and  29, we provide a detailed analysis of cross-dataset transfer performance across the MedMNIST collection for various self-supervised learning methods.
@@
[…]
@@
[…]
@@ E Conclusion
Novelty and Significance
Our work presents the most comprehensive analysis of self-supervised learning methods in medical imaging that systematically evaluates multiple critical aspects often overlooked, including OOD detection, cross-dataset generalizability, multi-domain pretraining, transfer learning and performance under label scarcity. Our finding that MoCo v3 consistently outperforms other methods—achieving superior performance in 5 out of 11 datasets while maintaining minimal accuracy degradation in cross-domain tasks and showing superior performance in OOD detection- provides a clear direction for method selection in medical applications. Equally significant is our discovery that ViT architectures, despite lower in-domain performance, demonstrate superior cross-dataset generalization and OOD detection capabilities—a critical consideration for real-world clinical deployment. Additionally, our observation that multi-modality training improves OOD detection while single-modality multi-domain training enhances accuracy offers practical guidance for designing robust medical imaging systems. These insights advance the understanding of how SSL techniques transfer across the unique challenges of medical domains and provide evidence-based recommendations for practitioners developing generalizable medical imaging AI.
F Key Findings and Empirical Observations
Our comprehensive analysis of self-supervised learning (SSL) methods in medical imaging revealed several important findings across different dimensions:
F.1 In-Domain Performance
  • Method Performance: MoCo v3 emerged as the most effective SSL method, achieving superior performance in 5 out of 11 datasets while maintaining minimal accuracy degradation (in 4 out of 11 datasets) for cross-domain generalization tasks.
  • Initialization Impact: ImageNet1K initialization generally improved in-domain performance across methods compared to random initialization. However, no clear superior approach emerged when comparing supervised ImageNet1K and self-supervised ImageNet1K weights for initializing self-supervised pretraining on medical images.
  • Architecture Comparison: ResNet architecture outperformed Vision Transformers (ViT) in the majority of in-domain evaluations. When trained from scratch, ResNet achieved higher accuracy in 8 out of 11 datasets. With pretrained weights, this advantage extended to 9 out of 11 datasets.
  • Initialization Effect on Architecture Gap: The performance gap between ResNet and ViT architectures narrowed when moving from random to ImageNet1K initialization, though the relative ranking remained largely unchanged (except for the Retina dataset).
  • Method-Specific Benefits: BYOL and DINO demonstrated remarkable improvement with ImageNet1K initialization, transitioning from among the worst performers to among the best performers.
  • Label Scarcity Effects: Under limited label availability, ViT exhibited larger accuracy drops compared to ResNet, a counterintuitive finding given its lower initial accuracy (where smaller drops would typically be expected due to diminishing returns).
F.2 Out-of-Distribution (OOD) Detection
  • Method Effectiveness: MoCo demonstrated superior OOD detection capabilities when trained from scratch.
  • Architecture Advantage: ViT consistently outperformed ResNet in OOD detection tasks across different methods and datasets.
  • Initialization Impact: The effect of ImageNet1K initialization on OOD detection performance was inconsistent, with significant variation across methods and datasets.
  • Architectural Trends: The ViT superiority trend generally persisted with ImageNet1K initialization, though with method-specific variations in relative performance.
F.3 Generalizability
  • Initialization Effect: ImageNet1K initialization improved cross-dataset generalizability, producing both better in-domain performance and more transferable representations.
  • Architecture Comparison: While ResNet-50 achieved higher in-domain performance than ViT-Small in most datasets (8/11 when trained from scratch and 9/11 when starting from ImageNet weights), ViT-Small showed different behavior in cross-dataset generalization. When pre-trained with ImageNet1K weights and evaluated on generalization tasks, ViT-Small reversed the in-domain trend by outperforming ResNet-50 in 8 out of 11 datasets, as shown in Figure 8. This suggests that ViT-Small benefits more from transfer learning in cross-dataset scenarios, despite its weaker in-domain performance.
  • Convergence with Pretraining: With pretrained weights, performance differences between architectures consistently approached zero, suggesting that pretraining reduces architecture-specific advantages.
F.4 Multi-Domain Effects
  • Single-Modality Impact: Single-modality multi-domain training (Organ{ACS}) generally improved in-domain accuracy but reduced OOD detection performance.
  • Multi-Modality Impact: Multi-modality training (Organ{AS}PnePath) typically decreased in-domain accuracy but consistently improved OOD detection capabilities.
These findings provide valuable insights into the complex interplay between self-supervised learning methods, architectural choices, initialization strategies, and dataset characteristics in medical imaging applications.